Minutes of the Geneva AMS group meeting on August 8, 2000

Geneva AMS group meetings

Next week there is a TIM( Technical Interchange Meeting and Safety Review Rehersal) meeting at CERN. Thursday morning there will be a progress report of all detectors. Electronics and GSI test beam meetings are also foreseen.

AMS-01 dismantling

Last thursday began finally the AMS01 tracker dismantling. The two external planes were taken out from the detector. The internal planes dismantling is foreseen for the end of this month.


PL shows the last kapton modifications with Wallraff holes for internal and external layers. Also he presented some offers from some companies. Meetings with them will be held next week at CERN.


Some advances on test system. Analog part design will be ready next week. The final card is a bit big (25x15cm) but it decouples everything. After that needed 1-2 weeks for PCB production and 1 week for assembly. Most likely will not be ready for September test beam. Anyway there is still a chance for use last days of test beam to test a baby ladder. Interface between PC and digital card is also possible. This week will be tested the velocity transfer between them.

Last week there were the first meeting for GSI electronic test beam. THis test beam will be between 3-10 november. Proposed to be there Giovanni(?), Pascal (army?) and Eduardo.

Rouland shows us the results about AD8051/2/4 op-amp total dose test. His presentation can be found here . In summary AD8051(TDR) and AD8052(hybrydK) shows good shape after radiation. Some problems with AD8054(hybridS). Next week will be repeated the test for AD8054 to be sure everything is correct.


Next week will be done the electrical prototype (12) for the test beam. No microsouldure before 21th august. Luckely test beam will begin 1st setptember instead of 25th august so we will be on time. After this one it will be constructed the baby ladder and lately the mechanical prototyopes.

On the 12-ladder mechanical prototype has been discovered a problem. The point is that the jig is 30um short probably due to a temperature variation in the clean room, before ~24C and now ~21C. Philipp is investigating. It seems it will be no important consequence for ladder assembly.

Future Meetings:

E. Cortina, Aug 8, 2000